### v1

# Benchmarking End-To-End Performance of AI-Based Chip Placement Algorithms

Abstract
The increasing complexity of modern very-large-scale integration (VLSI) design highlights the significance of Electronic Design Automation (EDA) technologies. Chip placement is a critical step in the EDA workflow, which positions chip modules on the canvas with the goal of optimizing performance, power, and area (PPA) metrics of final chip designs. Recent advances have demonstrated the great potential of AI-based algorithms in enhancing chip placement. However, due to the lengthy workflow of chip design, the evaluations of these algorithms often focus on intermediate surrogate metrics, which are easy to compute but frequently reveal a substantial misalignment with the end-to-end performance (i.e., the final design PPA). To address this challenge, we introduce ChiPBench, which can effectively facilitate research in chip placement within the AI community. ChiPBench is a comprehensive benchmark specifically designed to evaluate the effectiveness of existing AI-based chip placement algorithms in improving final design PPA metrics. Specifically, we have gathered $20$ circuits from various domains (e.g., CPU, GPU, and microcontrollers). These designs are compiled by executing the workflow from the verilog source code, which preserves necessary physical implementation kits, enabling evaluations for the placement algorithms on their impacts on the final design PPA. We executed six state-of-the-art AI-based chip placement algorithms on these designs and plugged the results of each single-point algorithm into the physical implementation workflow to obtain the final PPA results. Experimental results show that even if intermediate metric of a single-point algorithm is dominant, while the final PPA results are unsatisfactory. This suggests that the AI community should concentrate more on enhancing end-to-end performance rather than those intermediate surrogates. We believe that our benchmark will serve as an effective evaluation framework to bridge the gap between academia and industry.
1 Introduction
The exponential growth in the scale of integrated circuits (ICs), in accordance with Moore’s law, has posed significant challenges to chip design (Huang et al., 2021; Lopera et al., 2021). To handle the increasing complexity, many electronic design automation (EDA) tools have been developed to assist hardware engineers. As shown in Figure 1, EDA tools automate various steps in the chip design workflow, including high-level synthesis, logic synthesis, physical design, testing and verification (Huang et al., 2021; Sánchez et al., 2023).
Chip placement is a critical step in the chip design workflow, which aims to position chip modules on the canvas, with the goal of optimizing the performance, power, and area (PPA) metrics of final chip designs (Cheng et al., 2023; Shi et al., 2023; Lin et al., 2019). Traditionally, this is done manually by human expert designers, which costs much labor and necessitates much expert prior knowledge. Therefore, a lot of design automation methods, especially those AI-based algorithms, have been developed to automate this process. These methods mainly fall into two categories: optimization-based methods and reinforcement learning (RL)-based methods (Geng et al., 2024). Optimization-based methods employ traditional optimization algorithms, such as simulated annealing (SA) (Vashisht et al., 2020) and evolutionary algorithms (EA) (Shi et al., 2023) to directly address the large-scale optimization problem, exploring the design space to identify near-optimal solutions. In recent research, macro placement has been formulated as a Markov Decision Process (MDP), where the macro positions are determined sequentially (Mirhoseini et al., 2021). Reinforcement learning (RL) has emerged as a promising technique for this task due to its ability to continuously improve performance based on feedback from the environment through trial and error (Cheng & Yan, 2021; Cheng et al., 2022; Lai et al., 2022; 2023).
Figure 1: Illustration of the modern chip design workflow.
However, due to the lengthy workflow of chip design, the evaluations of these algorithms often focus on intermediate surrogate metrics, which are easy to compute but frequently reveal a substantial misalignment with the end-to-end performance (i.e., the final design PPA). On one hand, obtaining the end-to-end performance of a given chip placement solution requires a large amount of engineering efforts due to the lengthy workflow of chip design. In particular, we found that directly using the existing open-source EDA tools and chip placement datasets even usually fails to obtain the end-to-end performance. Thus, existing AI-based chip placement algorithms (Lai et al., 2023; Geng et al., 2024) train and evaluate learned models using intermediate surrogate metrics which are simple and easy to obtain. On the other hand, as the PPA metrics are reflected by many aspects that have not been adequately considered in the previous stages, there exhibits a critical gap between the surrogate metrics and the final PPA objectives. Therefore, this gap significantly limits the use of existing AI-based placement algorithms in practical industrial scenarios.
To address this challenge, we propose ChiPBench, a comprehensive benchmark specifically designed to evaluate the effectiveness of existing AI-based chip placement algorithms in improving final design PPA metrics. Appealing features of ChiPBench include its fully open source and reproducible characteristics, covering the entire EDA workflow from the source verilog code, and unifying the evaluation framework of AI-based chip placement methods using end-to-end performance. Thus, ChiPBench can effectively facilitate research in chip placement within the AI community by taking the first step towards a fully reproducible unified evaluation framework using the end-to-end performance. In terms of the dataset, we have gathered $20$ circuits from various domains (e.g., CPU, GPU, and microcontrollers). Then, these designs are compiled by executing the workflow from the verilog source code, which preserves sufficient physical implementation kits, enabling evaluations for the placement algorithms on their impacts on the final design PPA. In terms of the evaluated algorithms, we executed six state-of-the-art AI-based chip placement algorithms on the aforementioned designs and plugged the results of each single-point algorithm into the physical implementation workflow to obtain the final PPA results.
Experimental results show that even if intermediate metric of a single-point algorithm is dominant, while the final PPA results are unsatisfactory. Moreover, visualization experiments demonstrate that intermediate metrics have weak correlation with the final design PPA, emphasizing the importance of developing algorithms towards optimizing the final design PPA rather than intermediate metrics. This suggests that the AI community should concentrate more on enhancing end-to-end performance rather than those intermediate metrics. We believe that our benchmark will serve as an effective evaluation framework to bridge the gap between academia and industry.
We summarize our major contributions as follows. (1) Our proposed ChiPBench is a reproducible and unified evaluation framework of existing AI-based chip placement algorithms using end-to-end performance with fully open source EDA tools. This can effectively facilitate research in chip placement within the AI community. (2) We collected $20$ circuits from various domains, and construct a dataset by executing the EDA workflow from the verilog source code, preserving sufficient physical implementation kits for end-to-end performance evaluation. (3) We evaluate six state-of-the-art AI-based chip placement algorithms, including most popular AI-based chip placement algorithms. (4) Experiments demonstrate that intermediate metrics have weak correlation with the final design PPA, emphasizing the importance of developing algorithms towards optimizing the final design PPA rather than intermediate metrics.
2 Related Work
Datasets Some well-known EDA conferences, such as ISPD and ICCAD, host contests addressing EDA challenges and offer benchmarks with processed data for researchers. However, in the early years (e.g., ISPD2005 (Nam et al., 2005) and ICCAD2004 (Adya et al., 2009)), the provided datasets used overly simplified Bookshelf formats, which are abstracted versions of the actual design kits. Therefore, we cannot evaluate the final PPA of the placement results on those datasets. Recently, ISPD2015 (Bustany et al., 2015) and ICCAD2015 (Kim et al., 2015) have offered benchmarks and datasets closer to real-world applications, including necessary netlist, library, and design exchange files, broadening their utility slightly. Nevertheless, they still lack the essential information (e.g., necessary design kits) to run the open-source EDA tools such as OpenROAD (Kahng & Spyrou, 2021). Beyond these conferences, some other datasets have been developed in various directions. For example, the EPFL (Amarú et al., 2015) benchmarks and the larger OpenABC-D Chowdhury et al. (2021) dataset concentrated on synthetic netlists, primarily for testing modern logic optimization tools with a focus on logic synthesis. CircuitNet 2.0 Jiang et al. (2023), on the other hand, shifted the focus towards providing multi-modal data for prediction tasks, enhancing the capability for various prediction tasks through the use of diverse data modalities. Compared with previous efforts, our proposed dataset focuses on the entire EDA workflow. It provides complete files for each case and necessary design kits, such as timing constraints, library files, and LEF files, offering a comprehensive dataset that supports all stages of design and fosters a more integrated approach to chip design and evaluation.
Placement Algorithms Recent advancements in AI technology within the EDA field have led to a variety of AI-based chip placement algorithms. (1) Black-Box Optimization methods. Simulated Annealing (Cheng et al., 2023) provides a probabilistic method for finding a good approximation of the global optimum. Wire-Mask-Guided Black-Box Optimization (Shi et al., 2023) uses a wire-mask-guided greedy procedure to optimize macro placement efficiently. (2) Analytical methods. DREAMPlace (Lin et al., 2019) uses deep learning toolkits to achieve over a 30x speedup in placement tasks. AutoDMP (Agnesina et al., 2023) leverages DREAMPlace for the concurrent placement of macros and standard cells, enhancing macro placement quality. (3) Reinforcement Learning methods. MaskPlace (Lai et al., 2022) treats chip placement as a visual representation learning problem, reducing wirelength and ensuring zero overlaps. ChiPFormer (Lai et al., 2023) employs offline reinforcement learning, fine-tuning on unseen chips for better efficiency. The evaluation of these algorithms mainly focuses on intermediate metrics. In contrast, we utilized the end-to-end performance to evaluate six existing AI-based chip placement algorithms, encompassing a significant portion of mainstream AI-based placement algorithms.
3 Background on Electronic Design Automation
Electronic Design Automation (EDA) is a suite of software tools vital for designing and developing electronic systems, primarily integrated circuits (ICs). These tools enable electrical engineers to efficiently transform innovative concepts into functional products, addressing the increasing complexity and demands of modern chip design. EDA optimizes the entire design process from schematic capture to layout and fabrication, reducing time-to-market and enhancing design precision and sophistication. In the chip design workflow, EDA tools support various functions: they perform simulations to verify circuit behavior, execute synthesis to convert high-level descriptions to gate-level implementations, and manage physical layouts to ensure designs can be realized in silicon. As shown in Figure 1, the EDA design flow includes several key stages: logic synthesis  floorplanning, placement, Clock Tree Synthesis (CTS), and routing. Below are concise descriptions of each stage, illustrating their importance in the integrated circuit development process.
Logic Synthesis transforms a high-level circuit description into an optimized gate-level netlist implementation (Berndt et al., 2022; Wang et al., 2024a; b). Floorplan involves deciding the layout of major components within an integrated circuit. During this stage, large blocks and core components are positioned in a way that balances signal integrity, power distribution, and area utilization. Placement involves assigning specific locations to various circuit components within the core area of the chip, following the floorplanning stage. The primary objective of this stage is to strategically place the components to optimize performance metrics such as delay and power consumption while ensuring adherence to design rules (Geng et al., 2024). Clock Tree Synthesis(CTS) creates a clock distribution network within an IC to minimize those clock effects, and ensure the correct timing synchronization for circuit operation. Routing involves creating the physical paths for electrical connectivity between various components on the IC as per the netlist. This stage must handle multiple layers of the chip, avoid obstacles, manage signal integrity, and meet all electrical and timing constraints (Cheng et al., 2022).
Chip Placement The placement stage is crucially divided into two distinct phases: macro placement and cell placement. (1) Macro placement is a critical very large-scale integration (VLSI) physical design problem that targets the arrangement of larger components, such as SRAMs and clock generators—often called macros. This phase significantly impacts the chip’s overall floorplan and essential design parameters like wirelength, power, and area. (2) Following this, the standard cell placement phase addresses the arrangement of the more numerous and smaller standard cells, which serve as the fundamental building blocks of digital designs. This phase typically utilizes analytical solvers to secure an optimized configuration that not only minimizes wirelength but also enhances the electrical performance of the chip.
4 Dataset
4.1 Description of Designs
Due to the oversimplification of datasets in early years, there exists a significant gap between these datasets and real-world applications. For instance, the usually used Bookshelf format (Nam et al., 2005; Adya et al., 2009) is overly simplified so that placement results given in such format are inapplicable for the subsequent stages to obtain a valid final design. Some later datasets (Kim et al., 2015) provide the LEF/DEF and necessary files for running these stages, but the contained circuits are still limited and they still lack some information for open-source tools like OpenROAD to work. For instance, the library file lacks buffer definitions, which is necessary for the clock tree synthesis phase, and the LEF file has incomplete layer definitions, which hinders the routing phase. To address this issue, we construct a dataset with comprehensive physical implementation information across the entire flow. Our dataset involves collecting a series of designs spanning various domains, including components such as CPUs, GPUs, network interfaces, image processing technologies, IoT devices, cryptographic units, and microcontrollers. Additionally, the dataset features a diverse array of sizes, with cell ranges from thousands to nearly a million. The statistics for each case is detailed in Table 1, and we defer more details to Table 4 in Appendix C.
Table 1: Statistics of designs used in our benchmark.
Id Design #Cells #Nets #Macros #Pins #IOs #Edges
1 CAN-Bus (Tommydag, ) 815 935 0 2637 13 935
2 iot shield (brmcfarl, ) 904 1006 0 2995 33 974
3 FPGA-CAN (WangXuan95, ) 140848 178913 0 532024 4 176472
4 subrisc (Hara-Laboratory, ) 859382 1103295 0 3359066 34 1092653
5 sha256 (secworks, ) 10120 12283 0 38758 77 12176
6 8051 (lajanugen, ) 13865 16424 0 50848 10 16174
7 ariane133 (The-OpenROAD-Project, ) 168551 184856 132 592261 495 183142
8 bp (The-OpenROAD-Project, ) 301030 333364 24 984093 1198 333364
9 bp_be (The-OpenROAD-Project, ) 50881 58428 10 182949 3029 58092
10 camera interface (AngeloJacobo, ) 332 340 0 979 29 316
11 image processing (suntodai, ) 2333 3270 0 7823 64 3170
12 picorv (YosysHQ, ) 8851 10531 0 32195 409 10470
13 serv (olofk, ) 1291 1482 0 3915 306 1403
14 ariane136 (The-OpenROAD-Project, ) 175248 191081 136 609834 495 187911
15 bp_fe (The-OpenROAD-Project, ) 33206 36379 11 111510 2511 36203
16 chameleon (The-OpenROAD-Project, ) 61192 3396 4 10539 145 3285
17 riscv32i (The-OpenROAD-Project, ) 12740 10947 6 35370 135 9875
18 swerv_wrapper (The-OpenROAD-Project, ) 96435 105026 28 354652 1416 104565
19 toygpu (openrisc, ) 368081 466513 0 1399167 11 461675
20 mor1kx (matt-kimball, ) 104293 130743 0 374983 576 125979
4.2 Dataset Generation Pipeline
We use OpenROAD (Kahng & Spyrou, 2021), an open-source EDA tool, for generating our dataset. OpenROAD integrates various tools, such as Yosys (Wolf, 2016) for logic synthesis, TritonMacroPlacer for macro placement, RePlAce (Cheng et al., 2018) for cell placement, TritonCTS for clock tree synthesis, and TritonRoute for detailed routing. The choice of open-source tool allows for full reproducibility of our results and supports the promotion of the open-source community, ensuring that all generated data and methodologies are open-source. The initial dataset generation starts with Verilog files as raw data. OpenROAD performs logical synthesis to convert these high-level descriptions into a netlist, detailing the electrical connections among circuit components. This netlist is then used by OpenROAD’s integrated floorplanning tool to configure the physical layout of the circuit on silicon. The resulting design from the floorplanning stage is converted into LEF/DEF files by OpenROAD, facilitating the application of subsequent placement algorithms. Simultaneously, we complete the EDA design flow through OpenROAD, generating data at subsequent stages, including placement, CTS, and routing.
5 Algorithms
AI-based chip placement algorithms can be roughly grouped into three categories: black-box optimization (BBO) methods, analytical methods (gradient-based methods), and reinforcement learning (RL) methods. Each category frames the placement task as an optimization problem but adopts distinct objectives and methodologies. We present details as follows.
5.1 Black-Box-Optimization (BBO) Methods
A straightforward intuition is to view the chip placement task as a black-box-optimization (BBO) problem, where the inner workings of the objective functions are inaccessible, and solutions are evaluated only based on the output metrics.
Simulated Annealing (SA) is a heuristic BBO optimization algorithm favored for its simplicity in implementation. Specifically, the SA algorithm generates solutions by perturbing the solution space and then assessing the resulting representation. Different methods have been developed to effectively map representations to placement solutions (Kirkpatrick et al., 1983; Sherwani, 2012; Ho et al., 2004; Shunmugathammal et al., 2020; Vashisht et al., 2020), such as sequence pair (Murata et al., 1996) and B-tree (Chang et al., 2000). Solutions are probabilistically accepted based on an annealing temperature to escape local optima in pursuit of a global optimum. Due to its simplicity in implementation, the SA algorithm often serves as a strong baseline in previous studies. In this work, we incorporate a specific SA implementation (Cheng et al., 2023) utilizing operations like swaps, shifts, and shuffles, and a cost function that balances wirelength, density, and congestion.
WireMask-EA (Shi et al., 2023) is a BBO framework recently appearing on the AI conference NeurIPS’2023. It leverages wiremask to greedily guide the mapping from genotypes to phenotypes. Here the concept of wiremask was first proposed by Lai et al. (2022), and it is a matrix representing the potential increases in HPWL when placing the next macro on the canvas. Building on this, Shi et al. (2023) implement different kinds of BBO algorithms, including random search (RS), evolutionary algorithm (EA), and Bayesian optimization (BO).
5.2 Analytical (Gradient-Based) Methods
Analytical methods formulate the optimization objective as an analytical function of module coordinates. This formulation enables efficient solutions through techniques like quadratic programming (Kahng et al., 2005; Viswanathan et al., 2007a; b; Spindler et al., 2008; Chen et al., 2008; Kim et al., 2012; Kim & Markov, 2012) and direct gradient descent (Lu et al., 2014; Cheng et al., 2018; Lin et al., 2019; 2020; Gu et al., 2020; Liao et al., 2022). This work focuses on the gradient-based algorithms, which are by far the more mainstream algorithms.
DREAMPlace (Liao et al., 2022) is a GPU-accelerated framework that leverages differentiable proxies, such as approximate HPWL, as optimization objectives. It was built upon the previous analytical placement algorithms, ePlace (Lu et al., 2014) and RePlAce (Cheng et al., 2018), yet significantly speeding up the placement process by using GPUs for acceleration. The series of versions of DREAMPlace introduces diverse differentiable proxies to better align the PPA improvement.
AutoDMP (Agnesina et al., 2023) extends DREAMPlace by automating hyperparameter tuning through multi-objective Bayesian optimization. It further speeds up the optimization process and reduces the heavy manual tuning efforts. At that time, this work showcased the promising potential of integrating GPU-accelerated algorithms with machine learning techniques for the automation of VLSI design.
5.3 Reinforcement Learning (RL) Methods
As VLSI systems grow in complexity, RL methods are being explored to enhance placement quality. GraphPlace (Mirhoseini et al., 2021) first models macro placement as a RL problem. Subsequently, DeepPR (Cheng & Yan, 2021) and PRNet (Cheng et al., 2022) establish a streamlined pipeline encompassing macro placement, cell placement, and routing. However, they treat density as a soft constraint, which may violate non-overlap constraint during training. Therefore, in this work, we mainly focus on MaskPlace and ChiPFormer, which are recent SOTA algorithms with hard non-overlapping constraints.
MaskPlace (Lai et al., 2022) represents the chip states as pixel-level visual inputs, including a wiremask (recording the HPWL increment for each grid), the viewmask (a global observation of the canvas), and the positionmask (to ensure non-overlapping constraint). Furthermore, it uses dense reward to boost the sample efficiency.
ChiPFormer (Lai et al., 2023) represents the first offline RL method. It is pretrained on various chips via offline RL and then fine-tuned on unseen chips for better efficiency. As a result, the time for placement is significantly reduced.
6 Evaluation
6.1 Evaluation Metrics
6.1.1 Final Design PPA Metrics
The primary goal of the entire Electronic Design Automation (EDA) workflow is to optimize the final PPA metrics. PPA stands for performance, power, and area—three crucial dimensions used to evaluate the quality of a chip product. These dimensions are assessed using several critical metrics, including worst negative slack (WNS), total negative slack (TNS), number of violating paths (NVP), power, and area. Optimizing these PPA metrics has been a major focus in the industry, approached through expert-designed heuristics. However, the challenge of PPA optimization has not been fully recognized within the AI community. Bridging this gap and improving the incorporation of AI strategies into PPA optimization are key goals of this benchmark.
In terms of specific metrics, Worst Negative Slack (WNS) and Total Negative Slack (TNS) are essential for assessing the timing performance of a chip circuit. Slack is the discrepancy between the expected and required arrival times of a signal, with negative slack indicating a timing violation. WNS pinpoints the most severe negative slack within a circuit, thus identifying the most critical timing issue. Conversely, TNS aggregates all negative slacks, providing a comprehensive view of the circuit’s overall timing challenges. Moreover, the number of violating paths (NVP) counts the paths that fail to meet the timing constraints, further illustrating the timing performance issues.
6.1.2 Intermediate Metrics
Commonly used intermediate surrogate metrics include Congestion, Wire Length (WL), Half Perimeter Wire Length (HPWL), and Macro HPWL (mHPWL). Congestion evaluates the density of wires in different chip regions. High congestion in certain areas can pose substantial challenges during the routing stage. While not a direct component of the PPA metrics, managing congestion effectively is essential to ensure that the chip can be successfully manufactured. Therefore, it is also considered as an evaluation metric in this paper. Congestion is typically estimated after the Clock Tree Synthesis (CTS) stages but before the detailed routing stage, allowing for adjusting macro placement and routing strategies to mitigate potential issues.
Wire Length (WL) is the total length of all wires connecting all modules in a chip. Half Perimeter Wire Length (HPWL) is the sum of half perimeters of bounding boxes that encompass all pins in each net. It is widely used as an estimation of WL and is obtained after cell placement. Macro HPWL (mHPWL) further simplifies HPWL by only considering the macros. It is favored in recent studies as it can be immediately obtained after macro placement. These metrics are thought to correlate with the final PPA, but they do not directly reflect the chip quality.
6.2 End-to-End Evaluation Workflow
We present an end-to-end evaluation workflow utilizing OpenROAD-flow-scripts (Kahng & Spyrou, 2021) for the various stages of the EDA design flow, as illustrated in Figure 2. All tools used in this workflow are open-source, providing a significant advantage over other workflows that rely on commercial software. This workflow is designed to offer a comprehensive assessment of optimization algorithms at any stage of the design flow.
Figure 2: Illustration of our end-to-end evaluation workflow.
Our dataset comprises design kits needed for each stage of the physical design flow. For evaluating any stage-specific algorithm, the output file from the preceding stage serves as the input for the algorithm under evaluation. The algorithm processes this input to generate its output, which is subsequently plugged into the OpenROAD design flow. Ultimately, final performance metrics such as TNS, WNS, Area, and Power are reported, providing a comprehensive end-to-end performance assessment.This method offers a holistic set of metrics that can evaluate the optimization effects of any stage-specific algorithm on the final chip design, providing consistency of metrics and avoiding the limitations of overly simplified metrics confined to a single stage.It is particularly beneficial for the optimization and development of various algorithms, ensuring their improvements translate to practical enhancements in chip design and foster the development of more efficient and effective open-source EDA tools through a robust framework for testing and improvement.
6.3 Experimental Setup
We apply the aforementioned workflow to evaluate six macro placement algorithms: SA, WireMask-EA, MaskPlace, ChiPFormer, and the default algorithm in OpenROAD. As most of these methods only support the circuit data in a BookShelf format, while the circuits in our used dataset are in a standard LEF/DEF, we start by converting the LEF/DEF files from the floorplan stage of our dataset to BookShelf format to serve as the input for the placement algorithms. After finishing the macro placement stage, we use DREAMPlace to finish the placement of the standard cells. In this stage it conduct two different operations, coarse placement and legalization, with the coarse placement splitting the cells simultaneously using gradient descent algorithms and the legalization stage correct the positions of cells subject to the non-overlapping constraint. The resulting placement files in BookShelf format are then converted back to DEF format using Dreamplace and reintroduced into the original flow. Finally, we report the final metrics, obtaining end-to-end evaluation results. Additionally, we perform placement using OpenRoad’s native Place method and complete the entire flow to obtain the final metrics for comparison with other algorithms.
7 Results and Discussions
7.1 Main Results
We evaluate the AI-based chip placement algorithms, including SA, WireMask-EA, DREAMPlace, AutoDMP, MaskPlace, and ChiPFormer, using both intermediate metrics and end-to-end performance. The results for macro placement are in Table 2. ChiPFormer and WireMask-EA demonstrated a significant reduction in MacroHPWL compared to OpenROAD using TritonMacroPlacer. WireMask-EA achieved the best performance in terms of MacroHPWL. While these AI-based placement algorithms showed good performance on several intermediate metrics, they perform poorly in terms of the end-to-end metrics compared to OpenRoad, particularly in Power, TNS, and Area. This outcome revealed a significant gap between the originally optimized MacroHPWL intermediate metrics and the final design PPA.
Table 2: The evaluation results of AI-based macro placement algorithms. MacroHPWL, HPWL, Wirelength are intermediate metrics, and the other metrics evaluate the end-to-end performance.
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 0.698 1.115 1.165 1.170 1.058 1.092 1.096 0.865 1.007
SA 0.938 1.028 1.089 1.087 1.045 1.189 1.799 1.136 1.009
DREAMPlace 0.946 1.092 1.161 1.176 1.045 0.988 1.368 1.043 1.006
AutoDMP 0.762 0.950 0.981 0.981 1.021 0.986 1.089 0.973 1.005
MaskPlace 1.744 1.153 1.146 1.149 1.053 1.185 1.462 1.044 1.009
ChiPFormer 0.728 1.078 1.118 1.117 1.061 1.036 2.097 1.222 1.010
OpenROAD 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Table 3: The evaluation results of AI-based standard cell placement algorithms.
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 0.990 1.002 1.013 0.983 1.227 4.948 6.135 1.002
AutoDMP 1.159 1.141 1.160 1.011 1.170 1.259 1.208 1.006
OpenROAD 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Cell Placement As shown in Table 3, DREAMPlace achieved the best results in the intermediate metrics of HPWL, and performed well in terms of Power. However, OpenROAD achieved the best results in WNS, TNS, NVP, and area, further demonstrating the inconsistency between intermediate metrics and final PPA results.
7.2 Correlation Analysis
In this section, we conduct compute and discuss the correlation between the core optimization indicator MacroHPWL, used in existing placement algorithms, and the final chip performance metrics such as WNS, TNS, and wirelength, obtained through the OpenROAD process.
We use the Pearson correlation coefficient (Cohen et al., 2009) to evaluate the strength of linear correlation between pairs of metrics. The formula for calculating the Pearson correlation coefficient takes the form of
$r=\frac{\sum(X_{i}-\overline{X})(Y_{i}-\overline{Y})}{\sqrt{\sum(X_{i}-\overline{X})^{2}\sum(Y_{i}-\overline{Y})^{2}}},$ (1)
where $X_{i}$ and $Y_{i}$ are the observations, and $\overline{X}$ and $\overline{Y}$ are the respective means.
Figure 3: Correlation Between Metrics
The results are shown in Figure 3. To calculate the correlation, the signs of all values are adjusted so that for all metrics the lower indicates the better. The results show that MacroHPWL only has a weak correlation with the Wirelength, which indicates that existing algorithms that optimize MacroHPWL do not lead to an optimization on the Wirelength. In contrast, HPWL shows a very strong positive correlation with actual Wirelength, indicating that HPWL works as an effective surrogate for approximating the Wirelength.
(a) WL-TNS/WNS based on Method
(b) WL-TNS/WNS based on Case
Figure 4: Correlations Between Wirelength and TNS/WNS, In the visualizations, points that share the same color represent data from (a) same method or (b) same case, respectively.
In addition to Wirelength, the final PPA metrics of the chip are associated with WNS/TNS, congestion, area, and power. The analyses shows that the correlation between MacroHPWL and these metrics is weak, indicating that optimization of MacroHPWL has minimal impact on these performance indicators. Moreover, the results in Figure 4 show that Wirelength exhibits weak correlations with WNS and TNS as well. This implies that even if a single-point algorithm successfully optimizes metrics such as Wirelength, the ultimate physical implementation might only enhance one aspect of the PPA metrics and may not effectively optimize the other dimensions. Therefore, more appropriate intermediate metrics are needed to better correlate with the actual PPA objectives.
7.3 Discussion
Our benchmark comprises design kits essential for each stage of the physical design flow, including netlists, libraries, rules, and constraints needed during the physical implementation stage. This thorough inclusion allows for a convenient and detailed evaluation of algorithms at any specific stage of the physical design flow, enabling researchers and practitioners to test and compare the effectiveness of their solutions in a realistic, end-to-end environment. We call on the AI researchers to pay more attention on the “shift-left” challenge from the real-world industrial scenarios, keeping towards the mission of bridging the huge gap between academic research and industrial applications.
8 Conclusion
This paper presents a comprehensive dataset that spans the entire spectrum of the EDA design process and an end-to-end evaluation method, which we used to assess several placement algorithms: DREAMPlace, AutoDMP, MaskPlace, ChipFormer, Wiremask-EA, and SA. Our evaluation revealed inconsistencies between the metrics currently emphasized by mainstream placement algorithms and the final performance outcomes. These findings highlight the need for a new perspective in the development of placement algorithms.
Appendix A Technical details
A.1 Experimental Details
In our workflow, it starts with OpenROAD. The file format for physical design in OpenROAD is OpenDB database (.odb), which contains LEF/DEF information. OpenROAD can be used to convert ODB to LEF/DEF, or LEF/DEF to ODB. Initially, the ODB after macro placement in the OpenROAD flow is converted to LEF/DEF files, serving as inputs for other macro placement algorithms. Since most existing macro placement algorithms only support bookshelf format input, DREAMPlace’s code is used to convert LEF/DEF to bookshelf (.nodes, .pl, .nets) format. It is important to note that when DREAMPlace reads the DEF file, it treats the blockages in the BLOCKAGES part as virtual macros, so when using DREAMPlace’s PlaceDB in Python to read macro information, the virtual macros must be excluded.
For ChiPFormer, use the default model and perform 100 online iterations to fine-tune the macro layout.For MaskPlace, run 3000 epochs.Other settings use the defaults.
In the method of ChipFormer and MaskPlace, the experiments were run on an NVIDIA GeForce RTX 2080 Ti, taking one day for all cases.For the other algorithms, we used 32 CPUs (Intel(R) Xeon(R) CPU E5-2667 v4 @ 3.20GHz), with a total time expenditure of two days.
Next, the macro layout (.pl) files are obtained after running ChiPFormer, MaskPlace, WireMask-EA, and SA, and then DREAMPlace is used to write the macro layout into DEF files. Since DREAMPlace does not modify the blockages in the DEF file, if blockages are defined in the DEF file, additional script modifications may be required.
After obtaining the macro-placed DEF files, they are converted to ODB using OpenROAD, followed by Tapcell and Welltie insertion, PDN generation, IO place, global place, detail place, and subsequent CTS and routing.
The performance of OpenROAD is limited; in the future, other global placement tools and detailed placement tools can be used to run the entire workflow.
A.2 Encountered Errors
Due to certain parsing bugs in OpenROAD, when converting DEF files to ODB files, it is necessary to ensure that no object names in the DEF file contain ”/”, otherwise, there will be issues during timing tests (issues will occur when writing and reading SPEF file).
Since various algorithms are academic and cannot directly set the minimum spacing between macros (which can be set in OpenROAD), the optimization of macro positions by these algorithms may result in macros being placed too closely. When integrating back into OpenROAD, the following errors might occur:
OpenROAD’s limited capabilities in various stages also contribute to these errors. During PDN, ”Unable to repair all channels.” might be encountered. During the global place stage, global placement might diverge. During the detail place stage, detailed placement might fail (because the surrounding space of certain cells makes it impossible for OpenROAD to find space for adjustments).
The code of ChiPFormer has certain issues and needs modifications to be applied to other cases. DREAMPlace directly uses the mix-size method with LEF/DEF for running macros.
For the ICCAD 2015 dataset, it includes the lib file, LEF/DEF files, netlist file, and sdc file. However, the lib file lacks buffer definitions, preventing CTS. The lef files have incomplete layer definitions, hindering routing.Additionaly, OpenRoad does not support the syntax of the ICCAD 2015 sdc file.
Appendix B Limitation
Our dataset currently has limitations in terms of data volume. In the future, we aim to increase the dataset’s size and include more cases from various domains. This will enhance its generalizability and robustness, making it a more comprehensive and valuable resource for researchers . By expanding the dataset, we will improve its applicability across different areas of chip design and EDA, further refining and perfecting this important resource.
Appendix C More Results
The detailed descriptions of each design are in Table 4. All the raw data from the experiment are in Tables 5-22. Methods or designs not shown in the table failed to complete the entire process due to the errors discussed earlier.
Table 4: Detailed descriptions of our collected designs.
Id Design Description
1 CAN-BusTommydag A CAN bus Controller
2 iot shieldbrmcfarl IoT Shield for the Intel Galileo Development Board
3 FPGA-CANWangXuan95 A lightweight CAN bus controller
4 subriscHara-Laboratory Simple Instruction-Set Computer for IoT edge devices
5 sha256secworks Hardware implementation of the SHA-256 cryptographic hash function
6 8051lajanugen FPGA implementation of the 8051 Microcontroller
7 ariane133 The-OpenROAD-Project Ariane Core
8 black parrot The-OpenROAD-Project Full 64-bit RISC-V Core with Cache Coherence Directory
9 bp_be The-OpenROAD-Project Back-end of a 64-bit RISC-V Core with Cache Coherence Directory
10 camera interface AngeloJacobo Camera interface
11 image processing suntodai Image processing
12 picorv YosysHQ A Size-Optimized RISC-V CPU
13 serv olofk An award-winning bit-serial RISC-V core
14 ariane136 The-OpenROAD-Project Ariane Core
15 bp_fe The-OpenROAD-Project Front-end of a 64-bit RISC-V Core with Cache Coherence Directory
16 chameleon The-OpenROAD-Project SoC core
17 riscv32i The-OpenROAD-Project Simple RISCV 32I core
18 swerv_wrapper The-OpenROAD-Project SweRV RISC-V Core 1.1 from Western Digital
19 toygpu openrisc A simple GPU on a TinyFPGA BX
20 mor1kx matt-kimball an OpenRISC processor IP core
Table 5: The results of ariane133
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 1124169.895 5027796.1 0.227 6609025 0.369 -0.581 -687.560 2662 770528
SA 1683330.105 5109178.9 0.227 6594317 0.365 -0.812 -1368.340 3253 764028
DREAMPlace 1111023.474 4835183.2 0.222 6460013 0.370 -0.431 -578.455 2181 768426
AutoDMP 828592.368 4198935.5 0.195 5662040 0.351 -0.481 -622.147 2280 760842
MaskPlace 4444289.000 6258454.3 0.260 7720348 0.373 -0.847 -1553.500 3277 769790
ChiPFormer 1253799.737 4996747.8 0.229 6650789 0.372 -0.617 -933.311 2707 772172
OpenROAD 2685856.053 5226018.8 0.235 6817628 0.358 -0.492 -609.207 2279 770622
Table 6: The results of ariane136
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 1181404.421 5308713.2 0.243 7034760 0.578 -1.688 -4248.100 3546 791551
SA 1693986.737 5584824.9 0.252 7296870 0.574 -1.711 -4116.460 3662 792731
DREAMPlace 1067974.000 5157317.8 0.232 6885717 0.574 -1.380 -3293.090 3581 792608
AutoDMP 870515.895 4907336.9 0.229 6607637 0.569 -1.412 -3644.990 4071 791027
MaskPlace 4586233.421 6522052.1 0.290 8382804 0.586 -1.513 -3458.900 3608 796715
ChiPFormer 1269027.263 5078881.1 0.232 6705656 0.563 -1.348 -3266.360 3521 789667
OpenROAD 3067334.316 5540894.5 0.249 7176803 0.550 -1.318 -3204.870 3468 792446
Table 7: The results of black parrot
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 112560.500 7310237.2 0.422 9637354 0.252 -1.827 -2.222 23 815781
SA 151696.447 6548499.6 0.385 8775996 0.248 -1.896 -165.720 925 813755
AutoDMP 135090.737 6338160.8 0.369 8427105 0.247 -1.951 -60.856 605 811220
DREAMPlace 161581.000 6798714.6 0.382 8914932 0.244 -1.884 -198.213 1398 813703
MaskPlace 324368.447 6808053.5 0.393 8966592 0.249 -1.657 -102.182 987 814013
ChiPFormer 138943.842 6893462.7 0.394 8993077 0.250 -1.746 -463.346 1687 816983
OpenROAD 119123.132 6242894.2 0.360 8209505 0.246 -1.961 -71.882 862 810360
Table 8: The results of bpfe
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 45985.289 1566601.0 0.404 2206773 0.298 -1.013 -686.950 2263 234748
SA 59872.289 1289105.7 0.295 1633540 0.287 -1.060 -853.464 2459 237391
DREAMPlace 60280.000 1800046.2 0.498 2724715 0.301 -1.171 -808.973 2239 235832
AutoDMP 56107.000 1302271.3 0.293 1625619 0.292 -1.158 -765.161 2464 239894
MaskPlace 64024.000 1263243.8 0.285 1580587 0.285 -1.294 -641.347 1858 238130
ChiPFormer 45754.000 1282954.8 0.301 1670666 0.294 -0.967 -641.719 2457 235347
OpenROAD 53842.000 1167916.2 0.250 1389986 0.262 -1.007 -473.214 2441 231069
Table 9: The results of bpbe
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 66729.237 1826669.2 0.331 2288717.000 0.397 -1.829 -2427.590 4816 262891.000
SA 85984.158 1939541.5 0.410 2816347.000 0.411 -1.787 -3412.900 5612 264921.000
DREAMPlace 87304.667 1777230.3 0.317 2219014.333 0.386 -1.772 -2294.093 4885 262545.667
AutoDMP 86263.842 1713273.4 0.307 2124910.000 0.384 -1.819 -2436.980 4909 261516.000
MaskPlace 114314.553 1909505.5 0.345 2386914.000 0.398 -1.778 -3458.640 4730 261446.000
ChiPFormer 62168.447 2175102.7 0.411 2830643.000 0.417 -1.803 -2611.130 5127 266110.000
OpenROAD 76561.237 1940820.8 0.337 2332089.000 0.389 -1.827 -2514.450 5034 259572.000
Table 10: The results of swerv wrapper
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 474447.237 4186894.0 0.446 5209646 0.682 -1.127 -1256.030 1668 679048
SA 630919.500 4125523.0 0.455 5313519 0.682 -1.195 -880.450 1789 678647
DREAMPlace 865544.000 3673896.3 0.375 4631643 0.666 -0.889 -640.107 1165 675207
AutoDMP 450662.000 3171618.7 0.352 4130837 0.632 -0.711 -709.121 1495 672435
MaskPlace 1270726.500 4414508.8 0.461 5391952 0.681 -$1$743.094 1443 679852
ChiPFormer 437638.237 4056268.8 0.440 5144218 0.684 -1.079 -893.142 1756 679744
OpenROAD 666704.000 3187567.0 0.342 4015115 0.627 -0.978 -751.619 1531 673169
Table 11: The results of picorv32
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 88407.3 0.194 135544 0.048 -0.147 -0.521 29 18951.400
AutoDMP 132458.3 0.244 170435 0.050 -0.142 -0.525 37 19039.000
OpenROAD 87168.8 0.195 134211 0.048 -0.107 -0.661 38 18972.700
Table 12: The results of FPGA-CAN
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 1472841.9 0.181 2437446 0.738 -1.027 -678.865 17688 378643
AutoDMP 1437916.7 0.178 2393640 0.745 -0.892 -108.201 3864 379675
OpenROAD 1483498.8 0.180 2394544 0.742 -0.404 -30.331 1843 379193
Table 13: The results of CAN-Bus
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 3916.8 0.093 5683 0.005 -0.074 -0.074 1 1675.270
AutoDMP 4210.8 0.097 6024 0.005 -0.074 -0.074 1 1668.880
OpenROAD 3955.1 0.092 5623 0.005 -0.079 -0.079 1 1665.430
Table 14: The results of serv
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 7194.0 0.165 11128 0.007 -0.563 -10.333 81 2125.870
AutoDMP 10361.9 0.261 17455 0.008 -0.593 -11.893 120 2193.170
OpenROAD 7165.0 0.163 10638 0.007 -0.565 -10.210 73 2097.940
Table 15: The results of sha256
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 114376.4 0.203 170230 0.161 -0.542 -9.367 40 23476.400
AutoDMP 120832.7 0.206 173822 0.172 -0.573 -9.891 39 23398.700
OpenROAD 113841.8 0.202 168199 0.162 -0.528 -8.667 39 23426.900
Table 16: The results of iot_shield
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 5014.3 0.118 7346 0.006 -0.167 -2.457 27 1712.770
AutoDMP 5976.2 0.137 8653 0.006 -0.152 -2.367 29 1730.600
OpenROAD 5067.3 0.119 7349 0.006 -0.152 -2.186 20 1719.420
Table 17: The results of FPGA_image_processing
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 16556.5 0.123 23453 0.201 -0.924 -184.592 736 5797.470
AutoDMP 22203.0 0.160 30745 0.211 -0.930 -200.125 710 5812.630
OpenROAD 16689.0 0.120 22746 0.199 -0.919 -165.009 609 5758.100
Table 18: The results of 8051-master
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 142162.6 0.201 210127 0.075 -0.634 -16.460 35 29484.200
AutoDMP 131174.7 0.186 195536 0.074 -0.633 -15.643 35 29365.600
OpenROAD 141245.3 0.201 207737 0.073 -0.603 -16.664 41 29461.400
Table 19: The results of mor1kx
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 1723145.8 0.366 3421791 0.703 -1.863 -19564.300 26568 289570
AutoDMP 2258941.0 0.524 4893199 0.735 -3.879 -44135.900 26600 285455
OpenROAD 2006747.0 0.393 3661548 0.703 -0.796 -4928.920 25509 313065
Table 20: The results of toygpu
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 4347733.8 0.198 6784639 1.101 -1.503 -131.661 101 979293
OpenROAD 4657703.6 0.212 7185236 1.108 -1.294 -103.668 101 981911
Table 21: The results of subrisc
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 13920769.0 0.264 21051063 2.774 -0.847 -5742.680 29283 2341490
AutoDMP 13446106.8 0.238 19114559.0 2.697 -0.67 -150.178 466 2319150.0
OpenROAD 14386152.9 0.253 20108941 3.425 -0.782 -305.819 674 2334330.0
Table 22: The results of Camera_Interface
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 13920769.0 0.264 21051063 2.774 -0.847 -5742.680 29283 2341490
AutoDMP 13446106.8 0.238 19114559 2.697 -0.673 -150.178 466 2319150
OpenROAD 14386152.9 0.253 20108941 3.425 -0.782 -305.819 674 2334330
Appendix D License
The code and propose dataset will be publicly accessible. We include the following licenses for the raw data we used in this paper.