### v2

# Benchmarking End-To-End Performance of AI-Based Chip Placement Algorithms

Abstract
The increasing complexity of modern very-large-scale integration (VLSI) design highlights the significance of Electronic Design Automation (EDA) technologies. Chip placement is a critical step in the EDA workflow, which positions chip modules on the canvas with the goal of optimizing performance, power, and area (PPA) metrics of final chip designs. Recent advances have demonstrated the great potential of AI-based algorithms in enhancing chip placement. However, due to the lengthy workflow of chip design, the evaluations of these algorithms often focus on intermediate surrogate metrics, which are easy to compute but frequently reveal a substantial misalignment with the end-to-end performance (i.e., the final design PPA). To address this challenge, we introduce ChiPBench, which can effectively facilitate research in chip placement within the AI community. ChiPBench is a comprehensive benchmark specifically designed to evaluate the effectiveness of existing AI-based chip placement algorithms in improving final design PPA metrics. Specifically, we have gathered $20$ circuits from various domains (e.g., CPU, GPU, and microcontrollers). These designs are compiled by executing the workflow from the verilog source code, which preserves necessary physical implementation kits, enabling evaluations for the placement algorithms on their impacts on the final design PPA. We executed six state-of-the-art AI-based chip placement algorithms on these designs and plugged the results of each single-point algorithm into the physical implementation workflow to obtain the final PPA results. Experimental results show that even if intermediate metric of a single-point algorithm is dominant, while the final PPA results are unsatisfactory. This suggests that the AI community should concentrate more on enhancing end-to-end performance rather than those intermediate surrogates. We believe that our benchmark will serve as an effective evaluation framework to bridge the gap between academia and industry.
1 Introduction
The exponential growth in the scale of integrated circuits (ICs), in accordance with Moore’s law, has posed significant challenges to chip design (Huang et al., 2021; Lopera et al., 2021). To handle the increasing complexity, many electronic design automation (EDA) tools have been developed to assist hardware engineers. As shown in Figure 1, EDA tools automate various steps in the chip design workflow, including high-level synthesis, logic synthesis, physical design, testing and verification (Huang et al., 2021; Sánchez et al., 2023).
Chip placement is a critical step in the chip design workflow, which aims to position chip modules on the canvas, with the goal of optimizing the performance, power, and area (PPA) metrics of final chip designs (Cheng et al., 2023; Shi et al., 2023; Lin et al., 2019). Traditionally, this is done manually by human expert designers, which costs much labor and necessitates much expert prior knowledge. Therefore, a lot of design automation methods, especially those AI-based algorithms, have been developed to automate this process. These methods mainly fall into two categories: optimization-based methods and reinforcement learning (RL)-based methods (Geng et al., 2024). Optimization-based methods employ traditional optimization algorithms, such as simulated annealing (SA) (Vashisht et al., 2020) and evolutionary algorithms (EA) (Shi et al., 2023) to directly address the large-scale optimization problem, exploring the design space to identify near-optimal solutions. In recent research, macro placement has been formulated as a Markov Decision Process (MDP), where the macro positions are determined sequentially (Mirhoseini et al., 2021). Reinforcement learning (RL) has emerged as a promising technique for this task due to its ability to continuously improve performance based on feedback from the environment through trial and error (Cheng & Yan, 2021; Cheng et al., 2022; Lai et al., 2022; 2023).
Figure 1: Illustration of the modern chip design workflow.
However, due to the lengthy workflow of chip design, the evaluations of these algorithms often focus on intermediate surrogate metrics, which are easy to compute but frequently reveal a substantial misalignment with the end-to-end performance (i.e., the final design PPA). On one hand, obtaining the end-to-end performance of a given chip placement solution requires a large amount of engineering efforts due to the lengthy workflow of chip design. In particular, we found that directly using the existing open-source EDA tools and chip placement datasets even usually fails to obtain the end-to-end performance. Thus, existing AI-based chip placement algorithms (Lai et al., 2023; Geng et al., 2024) train and evaluate learned models using intermediate surrogate metrics which are simple and easy to obtain. On the other hand, as the PPA metrics are reflected by many aspects that have not been adequately considered in the previous stages, there exhibits a critical gap between the surrogate metrics and the final PPA objectives. Therefore, this gap significantly limits the use of existing AI-based placement algorithms in practical industrial scenarios.
To address this challenge, we propose ChiPBench, a comprehensive benchmark specifically designed to evaluate the effectiveness of existing AI-based chip placement algorithms in improving final design PPA metrics. Appealing features of ChiPBench include its fully open source and reproducible characteristics, covering the entire EDA workflow from the source verilog code, and unifying the evaluation framework of AI-based chip placement methods using end-to-end performance. Thus, ChiPBench can effectively facilitate research in chip placement within the AI community by taking the first step towards a fully reproducible unified evaluation framework using the end-to-end performance. In terms of the dataset, we have gathered $20$ circuits from various domains (e.g., CPU, GPU, and microcontrollers). Then, these designs are compiled by executing the workflow from the verilog source code, which preserves sufficient physical implementation kits, enabling evaluations for the placement algorithms on their impacts on the final design PPA. In terms of the evaluated algorithms, we executed six state-of-the-art AI-based chip placement algorithms on the aforementioned designs and plugged the results of each single-point algorithm into the physical implementation workflow to obtain the final PPA results.
Experimental results show that even if intermediate metric of a single-point algorithm is dominant, while the final PPA results are unsatisfactory. Moreover, visualization experiments demonstrate that intermediate metrics have weak correlation with the final design PPA, emphasizing the importance of developing algorithms towards optimizing the final design PPA rather than intermediate metrics. This suggests that the AI community should concentrate more on enhancing end-to-end performance rather than those intermediate metrics. We believe that our benchmark will serve as an effective evaluation framework to bridge the gap between academia and industry.
We summarize our major contributions as follows. (1) Our proposed ChiPBench is a reproducible and unified evaluation framework of existing AI-based chip placement algorithms using end-to-end performance with fully open source EDA tools. This can effectively facilitate research in chip placement within the AI community. (2) We collected $20$ circuits from various domains, and construct a dataset by executing the EDA workflow from the verilog source code, preserving sufficient physical implementation kits for end-to-end performance evaluation. (3) We evaluate six state-of-the-art AI-based chip placement algorithms, including most popular AI-based chip placement algorithms. (4) Experiments demonstrate that intermediate metrics have weak correlation with the final design PPA, emphasizing the critical importance of developing algorithms towards optimizing the final design PPA rather than intermediate metrics.
2 Related Work
Datasets Some well-known EDA conferences, such as ISPD and ICCAD, host contests addressing EDA challenges and offer benchmarks with processed data for researchers. However, in the early years (e.g., ISPD2005 (Nam et al., 2005) and ICCAD2004 (Adya et al., 2009)), the provided datasets used overly simplified Bookshelf formats, which are abstracted versions of the actual design kits. Therefore, we cannot evaluate the final PPA of the placement results on those datasets. Recently, ISPD2015 (Bustany et al., 2015) and ICCAD2015 (Kim et al., 2015) have offered benchmarks and datasets closer to real-world applications, including necessary netlist, library, and design exchange files, broadening their utility slightly. Nevertheless, they still lack the essential information (e.g., necessary design kits) to run the open-source EDA tools such as OpenROAD (Kahng & Spyrou, 2021). Beyond these conferences, some other datasets have been developed in various directions. For example, the EPFL (Amarú et al., 2015) benchmarks and the larger OpenABC-D Chowdhury et al. (2021) dataset concentrated on synthetic netlists, primarily for testing modern logic optimization tools with a focus on logic synthesis. CircuitNet 2.0 (Jiang et al., 2023), on the other hand, shifted the focus towards providing multi-modal data for prediction tasks, enhancing the capability for various prediction tasks through the use of diverse data modalities. Compared with previous efforts, our proposed dataset focuses on the entire EDA workflow. As shown in Table 1, a comparison between our dataset and mainstream datasets highlights the distinctions. It provides complete files for each case and necessary design kits, such as timing constraints, library files, and LEF files, offering a comprehensive dataset that supports all stages of physical implementation and fosters a more integrated approach to chip design and evaluation.
Table 1: Comparison Between Our Dataset and Existing Datasets
Dataset
Complete
Design Suite
Logic Synthesis
Support
Physical Design
Support
Full EDA Flow
Support
Large Scale
&Diversity
ISPD2005
ICCAD2015
EPFL Benchmarks
OpenABC-D
CircuitNet 2.0
ChiPBench (Ours)
Placement Algorithms Recent advancements in AI technology within the EDA field have led to a variety of AI-based chip placement algorithms. (1) Black-Box Optimization methods. Simulated Annealing (Cheng et al., 2023) provides a probabilistic method for finding a good approximation of the global optimum. Wire-Mask-Guided Black-Box Optimization (Shi et al., 2023) uses a wire-mask-guided greedy procedure to optimize macro placement efficiently. (2) Analytical methods. DREAMPlace (Lin et al., 2019) uses deep learning toolkits to achieve over a 30x speedup in placement tasks. AutoDMP (Agnesina et al., 2023) leverages DREAMPlace for the concurrent placement of macros and standard cells, enhancing macro placement quality. (3) Reinforcement Learning methods. MaskPlace (Lai et al., 2022) treats chip placement as a visual representation learning problem, reducing wirelength and ensuring zero overlaps. ChiPFormer (Lai et al., 2023) employs offline reinforcement learning, fine-tuning on unseen chips for better efficiency. The evaluation of these algorithms mainly focuses on intermediate metrics. In contrast, we utilized the end-to-end performance to evaluate six existing AI-based chip placement algorithms, encompassing a significant portion of mainstream AI-based placement algorithms.
3 Background on Electronic Design Automation
Electronic Design Automation (EDA) is a suite of software tools vital for designing and developing electronic systems, primarily integrated circuits (ICs). These tools enable electrical engineers to efficiently transform innovative concepts into functional products, addressing the increasing complexity and demands of modern chip design. EDA optimizes the entire design process from schematic capture to layout and fabrication, reducing time-to-market and enhancing design precision and sophistication. In the chip design workflow, EDA tools support various functions: they perform simulations to verify circuit behavior, execute synthesis to convert high-level descriptions to gate-level implementations, and manage physical layouts to ensure designs can be realized in silicon. As shown in Figure 1, the EDA design flow includes several key stages: logic synthesis, floorplanning, placement, Clock Tree Synthesis (CTS), and routing. Below are concise descriptions of each stage, illustrating their importance in the integrated circuit development process.
Logic Synthesis transforms a high-level circuit description into an optimized gate-level netlist (Berndt et al., 2022; Wang et al., 2024a; b). Floorplan involves deciding the layout of major components within an integrated circuit, positioning blocks and core components to balance signal integrity, power distribution, and area utilization. Placement involves assigning specific locations to various circuit components within the core area of the chip, following the floorplanning stage. The primary objective of this stage is to strategically place the components to optimize performance metrics such as delay and power consumption while ensuring adherence to design rules (Geng et al., 2024). Clock Tree Synthesis (CTS) creates a clock distribution network within an IC to minimize those clock effects, and ensure the correct timing synchronization for circuit operation. Routing involves creating the physical paths for electrical connectivity between various components on the IC as per the netlist. This stage must handle multiple layers of the chip, avoid obstacles, manage signal integrity, and meet all electrical and timing constraints (Cheng et al., 2022).
Chip Placement The placement stage is crucially divided into two distinct phases: macro placement and cell placement. (1) Macro placement is a critical very large-scale integration (VLSI) physical design problem that targets the arrangement of larger components, such as SRAMs and clock generators—often called macros. This phase significantly impacts the chip’s overall floorplan and essential design parameters like wirelength, power, and area. (2) Following this, the standard cell placement phase addresses the arrangement of the more numerous and smaller standard cells, which serve as the fundamental building blocks of digital designs. This phase typically utilizes analytical solvers to secure an optimized configuration that not only minimizes wirelength but also enhances the electrical and timing performance of the chip.
4 Dataset
4.1 Description of Designs
Due to the oversimplification of datasets in early years, there exists a significant gap between these datasets and real-world applications. For instance, the usually used Bookshelf format (Nam et al., 2005; Adya et al., 2009) is overly simplified so that placement results given in such format are inapplicable for the subsequent stages to obtain a valid final design. Some later datasets (Kim et al., 2015) provide the LEF/DEF and necessary files for running these stages, but the contained circuits are still limited and they still lack some information for open-source tools like OpenROAD to work. For instance, the library file lacks buffer definitions, which is necessary for the clock tree synthesis phase, and the LEF file has incomplete layer definitions, which hinders the routing phase. To address this issue, we construct a dataset with comprehensive physical implementation information across the entire flow. Our dataset involves collecting a series of designs spanning various domains, including components such as CPUs, GPUs, network interfaces, image processing technologies, IoT devices, cryptographic units, and microcontrollers. Additionally, the dataset features a diverse array of sizes, with cell ranges from thousands to nearly a million. Our dataset features a complete design suite that supports the full EDA flow and includes a diverse range of sizes and domains, as illustrated in Table 1. The statistics for each case is detailed in Table 2, and we defer more details to Table 7 in Appendix C.2.
Table 2: Statistics of designs used in our benchmark.
Id Design #Cells #Nets #Macros #Pins #IOs #Edges
1 8051 (lajanugen, ) 13865 16424 0 50848 10 16174
2 ariane136 (The-OpenROAD-Project, ) 175248 191081 136 609834 495 187911
3 ariane133 (The-OpenROAD-Project, ) 168551 184856 132 592261 495 183142
4 bp (The-OpenROAD-Project, ) 301030 333364 24 984093 1198 333364
5 bp_be (The-OpenROAD-Project, ) 50881 58428 10 182949 3029 58092
6 bp_fe (The-OpenROAD-Project, ) 33206 36379 11 111510 2511 36203
7 CAN-Bus (Tommydag, ) 815 935 0 2637 13 935
8 DE2_CCD_edge (suntodai, ) 2333 3270 0 7823 64 3170
9 dft48 Brendon Chetwynd, Kevin Bush, Kyle Ingols 48488 52575 68 125501 132 50654
10 FPGA-CAN (WangXuan95, ) 140848 178913 0 532024 4 176472
11 iot shield (brmcfarl, ) 904 1006 0 2995 33 974
12 mor1kx (openrisc, b) 104293 130743 0 374983 576 125979
13 or1200 openrisc (a) 43386 32195 20 97047 383 31958
14 OV7670_i2c (AngeloJacobo, ) 332 340 0 979 29 316
15 picorv (YosysHQ, ) 8851 10531 0 32195 409 10470
16 serv (olofk, ) 1291 1482 0 3915 306 1403
17 sha256 (secworks, ) 10120 12283 0 38758 77 12176
18 subrisc (Hara-Laboratory, ) 859382 1103295 0 3359066 34 1092653
19 swerv_wrapper (The-OpenROAD-Project, ) 96435 105026 28 354652 1416 104565
20 toygpu (matt-kimball, ) 368081 466513 0 1399167 11 461675
4.2 Dataset Generation Pipeline
We use OpenROAD (Kahng & Spyrou, 2021), an open-source EDA tool, for generating our dataset. OpenROAD integrates various tools, such as Yosys (Wolf, 2016) for logic synthesis, TritonMacroPlacer for macro placement, RePlAce (Cheng et al., 2018) for cell placement, TritonCTS for clock tree synthesis, and TritonRoute for detailed routing. The choice of open-source tool allows for full reproducibility of our results and supports the promotion of the open-source community, ensuring that all generated data and methodologies are open-source. The initial dataset generation starts with Verilog files as raw data. OpenROAD performs logical synthesis to convert these high-level descriptions into a netlist, detailing the electrical connections among circuit components. This netlist is then used by OpenROAD’s integrated floorplanning tool to configure the physical layout of the circuit on silicon. The resulting design from the floorplanning stage is converted into LEF/DEF files by OpenROAD, facilitating the application of subsequent placement algorithms. Simultaneously, we complete the EDA design flow through OpenROAD, generating data at subsequent stages, including placement, CTS, and routing.
5 Algorithms
AI-based chip placement algorithms can be roughly grouped into three categories: black-box optimization (BBO) methods, analytical methods (gradient-based methods), and reinforcement learning (RL) methods. Each category frames the placement task as an optimization problem but adopts distinct objectives and methodologies. We present details as follows.
5.1 Black-Box-Optimization (BBO) Methods
A straightforward intuition is to view the chip placement task as a black-box-optimization (BBO) problem, where the inner workings of the objective functions are inaccessible, and solutions are evaluated only based on the output metrics.
Simulated Annealing (SA) is a heuristic BBO optimization algorithm favored for its simplicity in implementation. Specifically, the SA algorithm generates solutions by perturbing the solution space and then assessing the resulting representation. Different methods have been developed to effectively map representations to placement solutions (Kirkpatrick et al., 1983; Sherwani, 2012; Ho et al., 2004; Shunmugathammal et al., 2020; Vashisht et al., 2020), such as sequence pair (Murata et al., 1996) and B-tree (Chang et al., 2000). Solutions are probabilistically accepted based on an annealing temperature to escape local optima in pursuit of a global optimum. Due to its simplicity in implementation, the SA algorithm often serves as a strong baseline in previous studies. In this work, we incorporate a specific SA implementation (Cheng et al., 2023) utilizing operations like swaps, shifts, and shuffles, and a cost function that balances wirelength, density, and congestion.
WireMask-EA (Shi et al., 2023) is a BBO framework that was recently introduced at the NeurIPS 2023 conference, positioning itself as an innovative approach in the intersection of AI and EDA. The framework utilizes a novel concept called wiremask, which plays a crucial role in guiding the mapping process from genotypes to phenotypes in a greedy manner. The wiremask concept was originally introduced by Lai et al. (2022), where it is defined as a matrix that predicts the potential increase in Half-Perimeter Wirelength (HPWL) for each subsequent macro placement on the design canvas. By estimating the wirelength increase, the wiremask helps in making informed decisions during the placement process, thereby potentially improving the quality of the layout. Building upon this concept, Shi et al. (2023) extended the framework to integrate several types of Black-Box Optimization (BBO) algorithms, including random search (RS), evolutionary algorithm (EA), and Bayesian optimization (BO), demonstrating the versatility of the approach in handling complex optimization tasks in chip design.
5.2 Analytical (Gradient-Based) Methods
Analytical methods formulate the optimization objective as an analytical function of module coordinates. This formulation enables efficient solutions through techniques like quadratic programming (Kahng et al., 2005; Viswanathan et al., 2007a; b; Spindler et al., 2008; Chen et al., 2008; Kim et al., 2012; Kim & Markov, 2012) and direct gradient descent (Lu et al., 2014; Cheng et al., 2018; Lin et al., 2019; 2020; Gu et al., 2020; Liao et al., 2022). This work focuses on the gradient-based algorithms, which are by far the more mainstream algorithms.
DREAMPlace (Liao et al., 2022) is a GPU-accelerated framework that leverages differentiable proxies, such as approximate HPWL, as optimization objectives. It was built upon the previous analytical placement algorithms, ePlace (Lu et al., 2014) and RePlAce (Cheng et al., 2018), yet significantly speeding up the placement process by using GPUs for acceleration. The series of versions of DREAMPlace introduces diverse differentiable proxies to better align the PPA improvement.
AutoDMP (Agnesina et al., 2023) extends DREAMPlace by automating hyperparameter tuning through multi-objective Bayesian optimization. It further accelerates the optimization process and reduces manual tuning efforts. At that time, this work showcased the promising potential of integrating GPU-accelerated algorithms with machine learning techniques for automating VLSI design.
5.3 Reinforcement Learning (RL) Methods
As VLSI systems grow in complexity, RL methods are being explored to enhance placement quality. GraphPlace (Mirhoseini et al., 2021) first models macro placement as a RL problem. Subsequently, DeepPR (Cheng & Yan, 2021) and PRNet (Cheng et al., 2022) establish a streamlined pipeline encompassing macro placement, cell placement, and routing. However, they treat density as a soft constraint, which may violate non-overlap constraint during training. Therefore, in this work, we mainly focus on MaskPlace and ChiPFormer, which are recent SOTA algorithms with hard non-overlapping constraints.
MaskPlace (Lai et al., 2022) represents the chip states as pixel-level visual inputs, including a wiremask (recording the HPWL increment for each grid), the viewmask (a global observation of the canvas), and the positionmask (to ensure non-overlapping constraint). Furthermore, it uses dense reward to boost the sample efficiency.
ChiPFormer (Lai et al., 2023) represents the first offline RL method. It is pretrained on various chips via offline RL and then fine-tuned on unseen chips for better efficiency. As a result, the time for placement is significantly reduced.
6 Evaluation
6.1 Evaluation Metrics
6.1.1 Final Design PPA Metrics
The primary goal of the entire Electronic Design Automation (EDA) workflow is to optimize the final PPA metrics. PPA stands for performance, power, and area—three crucial dimensions used to evaluate the quality of a chip product. These dimensions are assessed using several critical metrics, including worst negative slack (WNS), total negative slack (TNS), number of violating paths (NVP), power, and area. Optimizing these PPA metrics has been a major focus in the industry, approached through expert-designed heuristics. However, the challenge of PPA optimization has not been fully recognized within the AI community. Bridging this gap and improving the incorporation of AI strategies into PPA optimization are key goals of this benchmark.
In terms of specific metrics, Worst Negative Slack (WNS) and Total Negative Slack (TNS) are essential for assessing the timing performance of a chip circuit. Slack is the discrepancy between the expected and required arrival times of a signal, with negative slack indicating a timing violation. WNS pinpoints the most severe negative slack within a circuit, thus identifying the most critical timing issue. Conversely, TNS aggregates all negative slacks, providing a comprehensive view of the circuit’s overall timing challenges. Moreover, the number of violating paths (NVP) counts the paths that fail to meet the timing constraints, further illustrating the timing performance issues.
6.1.2 Intermediate Metrics
Commonly used intermediate surrogate metrics include Congestion, Wire Length (WL), Half Perimeter Wire Length (HPWL), and Macro HPWL (mHPWL). Congestion evaluates the density of wires in different chip regions. High congestion in certain areas can pose substantial challenges during the routing stage. While not a direct component of the PPA metrics, managing congestion effectively is essential to ensure that the chip can be successfully manufactured. Therefore, it is also considered as an evaluation metric in this paper. Congestion is typically estimated after the Clock Tree Synthesis (CTS) stages but before the detailed routing stage, allowing for adjusting macro placement and routing strategies to mitigate potential issues.
Wire Length (WL) is the total length of all wires connecting all modules in a chip. Half Perimeter Wire Length (HPWL) is the sum of half perimeters of bounding boxes that encompass all pins in each net. It is widely used as an estimation of WL and is obtained after cell placement. Macro HPWL (mHPWL) further simplifies HPWL by only considering the macros. It is favored in recent studies as it can be immediately obtained after macro placement. These metrics are thought to correlate with the final PPA, but they do not directly reflect the chip quality.
6.2 End-to-End Evaluation Workflow
Figure 2: Illustration of our end-to-end evaluation workflow.
We present an end-to-end evaluation workflow utilizing OpenROAD-flow-scripts (Kahng & Spyrou, 2021) for the various stages of the EDA design flow, as illustrated in Figure 2. All tools used in this workflow are open-source, providing a significant advantage over other workflows that rely on commercial software. This workflow is designed to offer a comprehensive assessment of optimization algorithms at any stage of the design flow.
Our dataset comprises design kits needed for each stage of the physical design flow. For evaluating any stage-specific algorithm, the output file from the preceding stage serves as the input for the algorithm under evaluation. The algorithm processes this input to generate its output, which is subsequently plugged into the OpenROAD design flow. Ultimately, final performance metrics such as TNS, WNS, Area, and Power are reported, providing a comprehensive end-to-end performance assessment.This method offers a holistic set of metrics that can evaluate the optimization effects of any stage-specific algorithm on the final chip design, providing consistency of metrics and avoiding the limitations of overly simplified metrics confined to a single stage.It is particularly beneficial for the optimization and development of various algorithms, ensuring their improvements translate to practical enhancements in chip design and foster the development of more efficient and effective open-source EDA tools through a robust framework for testing and improvement.
6.3 Experimental Setup
We apply the aforementioned workflow to evaluate six macro placement algorithms: SA, WireMask-EA, DREAMPlace, AutoDMP, MaskPlace, ChiPFormer, and the default algorithm in OpenROAD. Additionally, we also assess cell placement algorithms for DREAMPlace and AutoDMP. As most of these methods only support the circuit data in a BookShelf format, while the circuits in our used dataset are in a standard LEF/DEF, we start by converting the LEF/DEF files from the floorplan stage of our dataset to BookShelf format to serve as the input for the placement algorithms. After finishing the macro placement stage, the resulting placement files are then converted back to DEF format and reintroduced into the original flow. The resulting placement files in BookShelf format are then converted back to DEF and reintroduced into the original flow. Finally, we report the final metrics, obtaining end-to-end evaluation results. Additionally, we perform global placement and detailed placement using OpenROAD’s native Place method and complete the entire flow to obtain the final metrics for comparison with other algorithms. Our project is open-sourced on GitHub.
7 Results and Discussions
7.1 Main Results
Macro Placement We evaluate the AI-based chip placement algorithms, including SA, WireMask-EA, DREAMPlace, AutoDMP, MaskPlace, and ChiPFormer, using both intermediate metrics and end-to-end performance. The results for macro placement are in Table 3. ChiPFormer and WireMask-EA demonstrated a significant reduction in MacroHPWL compared to OpenROAD using TritonMacroPlacer. WireMask-EA achieved the best performance in terms of MacroHPWL. While these AI-based placement algorithms showed good performance on several intermediate metrics, they perform poorly in terms of the end-to-end metrics compared to OpenROAD, particularly in Power, TNS, and Area. This outcome revealed a significant gap between the originally optimized MacroHPWL intermediate metrics and the final design PPA.
Table 3: The evaluation results of AI-based macro placement algorithms. MacroHPWL, HPWL, Congestion are intermediate metrics, and the other metrics evaluate the end-to-end performance.
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 0.647 1.027 1.105 1.099 1.015 1.085 0.995 0.967 1.004
SA 0.836 1.044 1.099 1.097 1.062 1.121 1.311 1.109 1.013
DREAMPlace 0.857 0.974 1.049 1.059 1.015 1.112 1.025 1.038 0.999
AutoDMP 0.698 0.892 0.950 0.950 1.013 1.196 1.540 1.176 1.002
MaskPlace 1.681 1.119 1.148 1.148 1.051 1.014 0.978 0.903 1.014
ChiPFormer 0.681 0.976 1.027 1.024 1.015 1.031 1.355 1.223 0.981
OpenROAD 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Table 4: The evaluation results of AI-based standard cell placement algorithms.
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\downarrow$ TNS $\downarrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 0.981 0.999 1.008 0.987 1.321 4.678 5.313 0.996
AutoDMP 1.124 1.123 1.138 1.011 1.540 1.916 1.119 0.995
OpenROAD 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Cell Placement As shown in Table 4, DREAMPlace achieved the best results in the intermediate metrics of HPWL, and performed well in terms of Power. However, OpenROAD achieved the best results in WNS, TNS, and NVP, further demonstrating the inconsistency between intermediate metrics and final PPA results.
7.2 Correlation Analysis
In this section, we conduct compute and discuss the correlation between the core optimization indicator MacroHPWL, used in existing placement algorithms, and the final chip performance metrics such as WNS, TNS, and wirelength, obtained through the OpenROAD process.
We use the Pearson correlation coefficient (Cohen et al., 2009) to evaluate the strength of linear correlation between pairs of metrics. The formula for calculating the Pearson correlation coefficient takes the form of
$r=\frac{\sum(X_{i}-\overline{X})(Y_{i}-\overline{Y})}{\sqrt{\sum(X_{i}-\overline{X})^{2}\sum(Y_{i}-\overline{Y})^{2}}},$ (1)
where $X_{i}$ and $Y_{i}$ are the observations, and $\overline{X}$ and $\overline{Y}$ are the respective means.
The results are shown in Figure 4. To calculate the correlation, the signs of all values are adjusted so that for all metrics the lower indicates the better. The results show that MacroHPWL only has a weak correlation with the Wirelength, which indicates that existing algorithms that optimize MacroHPWL do not lead to an optimization on the Wirelength. In contrast, HPWL shows a very strong positive correlation with actual Wirelength, indicating that HPWL works as an effective surrogate for approximating the Wirelength.
(a) WL-TNS/WNS based on Method
(b) WL-TNS/WNS based on Case
Figure 3: Correlations Between Wirelength and TNS/WNS, In the visualizations, points that share the same color represent data from (a) same method or (b) same case, respectively.
Figure 4: Correlation Between Metrics
In addition to Wirelength, the final PPA metrics of the chip are associated with WNS/TNS, Area, and Power. The analyses shows that the correlation between MacroHPWL and these metrics is weak, indicating that optimization of MacroHPWL has minimal impact on these performance indicators. Moreover, the results in Figure 3 show that Wirelength exhibits weak correlations with WNS and TNS as well. This implies that even if a single-point algorithm successfully optimizes metrics such as Wirelength, the ultimate physical implementation might only enhance one aspect of the PPA metrics and may not effectively optimize the other dimensions. Therefore, more appropriate intermediate metrics are needed to better correlate with the actual PPA objectives.
7.3 Further Analysis
In this section, we analyze specific cases to assess the impact of different methods on PPA using ariane133 as the benchmark. Table 5 presents the experimental data, while Figure 5 shows the performance variations on the worst path caused by different placement algorithms. The AutoDMP method reduces wirelength but worsens timing, with smaller area and lower power. This is due to fewer buffers being added during timing repair, leading to degraded timing despite the reduced area and power. Moreover, wirelength and timing are not always proportional—critical paths still experience delays, indicating that global wirelength optimization may miss timing-sensitive paths, especially in worst-case scenarios. For more details, see Appendix C.1.
Table 5: The evaluation results of ariane133 under AI-based macro placement algorithms.
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 1124169 5065453.5 0.226 6583143 0.369 -0.417 -329.353 1970 349228
SA 1683330 5187015.8 0.230 6699484 0.365 -0.512 -650.399 2317 347043
MaskPlace 4444289 6253554.3 0.265 7853892 0.373 -0.349 -244.936 1636 357322
ChiPFormer 1253799 5019138.5 0.226 6581086 0.370 -0.553 -860.952 2703 349005
DREAMPlace 1111023 4826654.0 0.214 6348638 0.367 -0.540 -690.266 2307 348180
AutoDMP 828592 4250870.4 0.192 5694373 0.350 -0.982 -1913.660 3310 344091
OpenROAD 2685856 5260791.9 0.235 6825071 0.359 -0.498 -596.718 2274 352706
Figure 5: The images of the worst path for each method in ariane133.
7.4 Discussion
Our benchmark comprises design kits essential for each stage of the physical design flow, including netlists, libraries, rules, and constraints needed during the physical implementation stage. This thorough inclusion allows for a convenient and detailed evaluation of algorithms at specific stages of the physical design flow, enabling researchers and practitioners to test and compare the effectiveness of their solutions in a realistic, end-to-end environment. We call on the AI researchers to pay more attention on the “shift-left” challenge from the real-world industrial scenarios, keeping towards the mission of bridging the huge gap between academic research and industrial applications.
8 Conclusion
This paper presents a comprehensive dataset that spans the entire spectrum of the EDA design process and an end-to-end evaluation method, which we used to assess several placement algorithms: SA, WireMask-EA, DREAMPlace, AutoDMP, MaskPlace, and ChiPFormer. Our evaluation revealed inconsistencies between the metrics currently emphasized by mainstream placement algorithms and the final performance outcomes. These findings highlight the need for a new perspective in the development of placement algorithms.
Appendix A Technical details
A.1 Experimental Details
In our workflow, it starts with OpenROAD. The file format for physical design in OpenROAD is OpenDB database (.odb), which contains LEF/DEF information. OpenROAD can be used to convert ODB to LEF/DEF, or LEF/DEF to ODB. Initially, the ODB after macro placement in the OpenROAD flow is converted to LEF/DEF files, serving as inputs for other macro placement algorithms. Since most existing macro placement algorithms only support bookshelf format input, DREAMPlace’s code is used to convert LEF/DEF to bookshelf (.nodes, .pl, .nets) format. It is important to note that when DREAMPlace reads the DEF file, it treats the blockages in the BLOCKAGES part as virtual macros, so when using DREAMPlace’s PlaceDB in Python to read macro information, the virtual macros must be excluded.
For ChiPFormer, use the default model and perform 100 online iterations to fine-tune the macro layout.For MaskPlace, run 3000 epochs.Other settings use the defaults.
In the method of ChipFormer and MaskPlace, the experiments were run on an NVIDIA GeForce RTX 2080 Ti, taking one day for all cases.For the other algorithms, we used 32 CPUs (Intel(R) Xeon(R) CPU E5-2667 v4 @ 3.20GHz), with a total time expenditure of two days.
Next, the macro layout (.pl) files are obtained after running ChiPFormer, MaskPlace, WireMask-EA, and SA, and then DREAMPlace is used to write the macro layout into DEF files. Since DREAMPlace does not modify the blockages in the DEF file, if blockages are defined in the DEF file, additional script modifications may be required.
After obtaining the macro-placed DEF files, they are converted to ODB using OpenROAD, followed by Tapcell and Welltie insertion, PDN generation, IO place, global place, detail place, and subsequent CTS and routing.
The performance of OpenROAD is limited; in the future, other global placement tools and detailed placement tools can be used to run the entire workflow.
A.2 Encountered Errors
Due to certain parsing bugs in OpenROAD, when converting DEF files to ODB files, it is necessary to ensure that no object names in the DEF file contain ”/”, otherwise, there will be issues during timing tests (issues will occur when writing and reading SPEF file).
Since various algorithms are academic and cannot directly set the minimum spacing between macros (which can be set in OpenROAD), the optimization of macro positions by these algorithms may result in macros being placed too closely. When integrating back into OpenROAD, the following errors might occur:
OpenROAD’s limited capabilities in various stages also contribute to these errors. During PDN, ”Unable to repair all channels.” might be encountered. During the global place stage, global placement might diverge. During the detail place stage, detailed placement might fail (because the surrounding space of certain cells makes it impossible for OpenROAD to find space for adjustments).
The code of ChiPFormer has certain issues and needs modifications to be applied to other cases. DREAMPlace directly uses the mixed-size method with LEF/DEF for running macros.
For the ICCAD 2015 dataset, it includes the lib file, LEF/DEF files, netlist file, and sdc file. However, the lib file lacks buffer definitions, preventing CTS. The lef files have incomplete layer definitions, hindering routing.Additionally, OpenROAD does not support the syntax of the ICCAD 2015 sdc file.
Appendix B Limitation
Our dataset currently has limitations in terms of data volume. In the future, we aim to increase the dataset’s size and include more cases from various domains. This will enhance its generalizability and robustness, making it a more comprehensive and valuable resource for researchers . By expanding the dataset, we will improve its applicability across different areas of chip design and EDA, further refining and perfecting this important resource.
Appendix C More Results
C.1 Additional Analysis
In this section, we provide a detailed analysis of the layout results for the ariane133 benchmark, focusing on the observed relationship between wirelength, timing, area, and power. The AutoDMP method shows interesting yet seemingly contradictory outcomes—reduced wirelength with worse timing, smaller area, and lower power. To understand these results, we explore the factors that contribute to these trends and propose potential directions for placement algorithm improvements.
A key factor influencing timing in large-scale circuits is buffer insertion. In general, when the wirelength between two pins is large, signal transmission delay increases. Buffer insertion helps by breaking long interconnects into smaller segments, reducing the overall transmission delay. Furthermore, buffers enhance driving capacity by reducing the load on the signal driver, which helps lower signal delay and improve timing. However, fewer buffers result in smaller area and lower power but also lead to degraded timing performance. In the case of ariane133, the AutoDMP method inserts fewer buffers during timing repair, resulting in smaller overall area and power consumption. However, the lack of sufficient buffers leads to worse timing outcomes, as seen in the experimental data. This explains why, despite achieving minimized wirelength, the timing performance does not improve accordingly.
The relationship between wirelength and timing is often assumed to be directly proportional; however, the analysis of ariane133 demonstrates that this is not always the case. Although the total wirelength (as indicated by the HPWL) is reduced, the timing performance does not improve proportionally. In fact, congestion is lower and wirelength is minimized, but critical paths still suffer from significant delays. This suggests that global wirelength optimization does not always lead to improved timing, particularly for worst-case paths. Table 6 provides insights into this phenomenon by highlighting the inconsistencies between wirelength and timing slack on the worst paths. Specifically, the data reveals that longer wirelength paths do not always exhibit worse timing, indicating that the optimization strategy may not have sufficiently accounted for timing-sensitive paths.
To address these issues and improve overall placement quality, several optimization strategies can be explored. One potential solution to mitigate timing degradation is to apply weighted optimization that emphasizes critical paths. By giving more weight to timing-sensitive paths during wirelength reduction, it would be possible to achieve better alignment between wirelength optimization and timing improvement. Another approach is to adopt mixed-size placement techniques that optimize both macros and standard cells simultaneously. Since macro wirelength (macro HPWL) has a weaker correlation with backend performance metrics, this approach could lead to a more balanced placement outcome, particularly for designs where macro placement significantly impacts PPA. Additionally, machine learning models that predict PPA based on placement features could provide valuable insights during placement optimization. These models would allow the algorithm to dynamically adjust its strategy based on predicted PPA outcomes, leading to more informed trade-offs between PPA.
Table 6: Comparison of the worst path wirelength, delay, and slack for various methods on the ariane133 layout.
Method WireMask-EA SA MaskPlace ChiPFormer DREAMPlace AutoDMP OpenROAD
wirelength 3892 4107 3529 5761 4612 4318 4121
delay 4.36 4.57 4.25 4.49 4.54 4.92 4.44
slack -0.42 -0.54 -0.35 -0.55 -0.54 -0.98 -0.51
C.2 Raw Data
All the raw data from the experiment are in Tables 8-27.
Table 7: Detailed descriptions of our collected designs.
Id Design Description
1 8051lajanugen FPGA implementation of the 8051 Microcontroller
2 ariane133 The-OpenROAD-Project Ariane Core
3 ariane136 The-OpenROAD-Project Ariane Core
4 bp The-OpenROAD-Project Full 64-bit RISC-V Core with Cache Coherence Directory
5 bp_be The-OpenROAD-Project Back-end of a 64-bit RISC-V Core with Cache Coherence Directory
6 bp_fe The-OpenROAD-Project Front-end of a 64-bit RISC-V Core with Cache Coherence Directory
7 CAN-BusTommydag A CAN bus Controller
8 DE2_CCD_edge suntodai Image processing
9 dft48Brendon Chetwynd, Kevin Bush, Kyle Ingols DFT design
10 FPGA-CANWangXuan95 A lightweight CAN bus controller
11 iot shieldbrmcfarl IoT Shield for the Intel Galileo Development Board
12 mor1kx openrisc (b) an OpenRISC processor IP core
13 or1200openrisc (a) OpenRISC 1200 implementation
14 OV7670_i2c AngeloJacobo Camera interface
15 picorv YosysHQ A Size-Optimized RISC-V CPU
16 serv olofk An award-winning bit-serial RISC-V core
17 sha256secworks Hardware implementation of the SHA-256 cryptographic hash function
18 subriscHara-Laboratory Simple Instruction-Set Computer for IoT edge devices
19 swerv_wrapper The-OpenROAD-Project SweRV RISC-V Core 1.1 from Western Digital
20 toygpu matt-kimball A simple GPU on a TinyFPGA BX
Table 8: The results of ariane133
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 1124169 5065453.5 0.226 6583143 0.369 -0.417 -329.353 1970 349228
SA 1683330 5187015.8 0.230 6699484 0.365 -0.512 -650.399 2317 347043
MaskPlace 4444289 6253554.3 0.265 7853892 0.373 -0.349 -244.936 1636 357322
ChiPFormer 1253799 5019138.5 0.226 6581086 0.370 -0.553 -860.952 2703 349005
DREAMPlace 1111023 4826654.0 0.214 6348638 0.367 -0.540 -690.266 2307 348180
AutoDMP 828592 4250870.4 0.192 5694373 0.350 -0.982 -1913.660 3310 344091
OpenROAD 2685856 5260791.9 0.235 6825071 0.359 -0.498 -596.718 2274 352706
Table 9: The results of ariane136
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 1181404 5278467.4 0.239 6945252 0.570 -1.726 -4268.610 3628 358740
SA 1693986 5608216.9 0.248 7208805 0.571 -1.620 -3917.960 3487 359532
MaskPlace 4586233 6544926.4 0.285 8245218 0.583 -1.508 -3677.480 3966 363704
ChiPFormer 1269027 5244085.2 0.237 6869186 0.566 -1.385 -3603.530 3609 358773
DREAMPlace 1067974 5202240.5 0.231 6831531 0.571 -1.358 -3269.220 4350 356706
AutoDMP 870515 4993936.2 0.228 6590972 0.562 -1.553 -4236.020 4431 356714
OpenROAD 3067334 5561870.0 0.246 7124942 0.546 -1.420 -3707.930 3473 360120
Table 10: The results of bp
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 112560 7526021.6 0.437 10002159 0.254 -1.936 -21.812 326 490622
SA 151696 6927899.3 0.398 9113560 0.253 -1.965 -143.317 969 489850
MaskPlace 324368 7135790.7 0.407 9307133 0.253 -1.625 -44.217 478 489867
ChiPFormer 138943 6857073.6 0.392 8970666 0.250 -1.752 -502.977 2179 490305
DREAMPlace 161581 6989990.7 0.409 9347541 0.250 -2.108 -14.609 192 489693
AutoDMP 135090 6568910.6 0.376 8610145 0.247 -1.791 -254.786 1821 486445
OpenROAD 119123 6389902.6 0.365 8365647 0.247 -1.674 -219.229 1095 484494
Table 11: The results of bp_be
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 66729 2678100.1 0.524 3574875 0.467 -2.142 -4093.970 5131 123966
SA 85984 2568796.6 0.557 3788559 0.459 -2.494 -5510.140 5381 124667
MaskPlace 114314 2548230.4 0.513 3522611 0.456 -2.439 -4459.870 5173 124938
ChiPFormer 62168 2525703.6 0.524 3572070 0.425 -2.169 -3541.820 5100 110904
DREAMPlace 87304 2533359.4 0.516 3518916 0.458 -2.163 -3648.020 6026 125221
AutoDMP 86263 2539005.8 0.468 3227167 0.453 -1.948 -3351.120 5045 125471
OpenROAD 76561 2370091.6 0.409 2835542 0.409 -1.862 -2343.290 5013 118786
Table 12: The results of bp_fe
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 45985 1863651.1 0.510 2783740 0.309 -1.666 -777.420 2628 77648.300
SA 59872 1603778.9 0.404 2218668 0.310 -1.179 -1236.880 2990 81919.500
MaskPlace 64024 1540944.3 0.355 1965209 0.309 -1.318 -771.363 2582 77881.100
ChiPFormer 45754 1544281.0 0.375 2073376 0.299 -1.197 -1000.190 2714 72723.900
DREAMPlace 60279 1975681.6 0.508 2823861 0.294 -1.117 -473.261 1849 77427.800
AutoDMP 56107 1560612.6 0.365 2023847 0.312 -1.253 -1198.430 2626 81294.400
OpenROAD 53842 1468799.8 0.319 1770176 0.285 -1.092 -491.698 1870 75884.500
Table 13: The results of dft48
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 335246296 1039236.4 0.114 1330932 0.239 -0.510 -439.378 2035 81353.700
SA 339003863 1815564.2 0.198 2305190 0.304 -0.882 -422.568 1546 84170.100
MaskPlace 1409226442 2117408.1 0.223 2552954 0.258 -0.381 -98.916 548 83185.400
ChiPFormer 431188529 1094111.0 0.119 1366989 0.242 -0.561 -235.601 974 80770.400
DREAMPlace 267700115 800165.5 0.112 1308094 0.234 -1.034 -801.257 2274 79407.400
AutoDMP 270017585 774418.4 0.091 1051008 0.232 -0.453 -125.948 797 80201.100
OpenROAD 752297935 1981412.2 0.198 2267002 0.257 -0.660 -393.111 1411 84300.200
Table 14: The results of or1200
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 102261063 1196931.6 0.035 1432301 0.049 -1.417 -1530.180 1754 66547.600
SA 113496323 1200696.8 0.034 1420640 0.059 -1.294 -1613.410 2674 66937.300
MaskPlace 208736767 1166784.1 0.033 1350820 0.059 -1.505 -2198.370 2684 66953.500
ChiPFormer 95035903 1089714.1 0.030 1262132 0.057 -1.208 -1415.960 2669 66256.900
DREAMPlace 153596940 1099937.2 0.031 1288746 0.057 -1.151 -1380.980 2670 65981.000
AutoDMP 122473700 1058807.4 0.033 1373410 0.058 -2.025 -3119.280 2760 65756.800
OpenROAD 187077195 1223792.9 0.034 1405193 0.060 -1.278 -1595.730 2677 67276.500
Table 15: The results of swerv_wrapper
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 474447 3893655.8 0.414 4854661 0.666 -1.027 -873.506 1518 205627
SA 630919 3811960.2 0.368 4310596 0.648 -0.962 -854.737 1811 200280
MaskPlace 1270726 4193788.0 0.452 5286392 0.690 -1.251 -1306.460 1696 206481
ChiPFormer 437638 4060188.1 0.428 5019849 0.674 -1.189 -1282.240 2139 205775
DREAMPlace 865544 3575327.7 0.366 4525348 0.646 -1.061 -780.196 1608 203896
AutoDMP 450662 3213155.0 0.372 4354901 0.648 -1.094 -773.339 1529 200823
OpenROAD 666704 3177043.0 0.339 3997163 0.628 -1.127 -1163.820 1515 202853
Table 16: The results of 8051-master
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 142162.6 0.201 210127 0.075 -0.634 -16.460 35 29484.200
AutoDMP 131174.7 0.186 195536 0.074 -0.633 -15.643 35 29365.600
OpenROAD 141245.3 0.201 207737 0.073 -0.603 -16.664 41 29461.400
Table 17: The results of CAN-Bus
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 3916.8 0.093 5683 0.005 -0.074 -0.074 1 1675.270
AutoDMP 4210.8 0.097 6024 0.005 -0.074 -0.074 1 1668.880
OpenROAD 3955.1 0.092 5623 0.005 -0.079 -0.079 1 1665.430
Table 18: The results of DE2_CCD_edge
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 16556.5 0.123 23453 0.201 -0.924 -184.592 736 5797.470
AutoDMP 22203.0 0.160 30745 0.211 -0.930 -200.125 710 5812.630
OpenROAD 16689.0 0.120 22746 0.199 -0.919 -165.009 609 5758.100
Table 19: The results of FPGA-CAN
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 1472841.9 0.181 2437446 0.738 -1.027 -678.865 17688 378643
AutoDMP 1437916.7 0.178 2393640 0.745 -0.892 -108.201 3864 379675
OpenROAD 1483498.8 0.180 2394544 0.742 -0.404 -30.331 1843 379193
Table 20: The results of OV7670_i2c
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 733113.0 0.215 1164763 0.625 -0.579 -82.237 3154 157496
AutoDMP 701777.0 0.201 1091553 0.612 -0.561 -23.030 1053 155624
OpenROAD 723012.1 0.208 1127120 0.617 -0.466 -30.612 2327 156210
Table 21: The results of iot_shield
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 5014.3 0.118 7346 0.006 -0.167 -2.457 27 1712.770
AutoDMP 5976.2 0.137 8653 0.006 -0.152 -2.367 29 1730.600
OpenROAD 5067.3 0.119 7349 0.006 -0.152 -2.186 20 1719.420
Table 22: The results of mor1kx
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 1723145.8 0.366 3421791 0.703 -1.863 -19564.300 26568 289570
AutoDMP 2258941.0 0.524 4893199 0.735 -3.879 -44135.900 26600 285455
OpenROAD 2006747.0 0.393 3661548 0.703 -0.796 -4928.920 25509 313065
Table 23: The results of picorv32
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 88407.3 0.194 135544 0.048 -0.147 -0.521 29 18951.400
AutoDMP 132458.3 0.244 170435 0.050 -0.142 -0.525 37 19039.000
OpenROAD 87168.8 0.195 134211 0.048 -0.107 -0.661 38 18972.700
Table 24: The results of serv
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 7194.0 0.165 11128 0.007 -0.563 -10.333 81 2125.870
AutoDMP 10361.9 0.261 17455 0.008 -0.593 -11.893 120 2193.170
OpenROAD 7165.0 0.163 10638 0.007 -0.565 -10.210 73 2097.940
Table 25: The results of sha256
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 114376.4 0.203 170230 0.161 -0.542 -9.367 40 23476.400
AutoDMP 120832.7 0.206 173822 0.172 -0.573 -9.891 39 23398.700
OpenROAD 113841.8 0.202 168199 0.162 -0.528 -8.667 39 23426.900
Table 26: The results of subrisc
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 13920769.0 0.264 21051063 2.774 -0.847 -5742.680 29283 2341490
AutoDMP 13446106.8 0.238 19114559 2.697 -0.673 -150.178 466 2319150
OpenROAD 14386152.9 0.253 20108941 3.425 -0.782 -305.819 674 2334330
Table 27: The results of toygpu
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 4347733.8 0.198 6784639 1.101 -1.503 -131.661 101 979293
AutoDMP 4444860.0 0.194 6595166 1.109 -2.418 -201.797 107 967780
OpenROAD 4657703.6 0.212 7185236 1.108 -1.294 -103.668 101 981911
Appendix D License
The code and propose dataset will be publicly accessible. We include the following licenses for the raw data we used in this paper.